OpenStudy (anonymous):

Calculate the volume of the solid, bounded by surfaces: z=x^2+y^2; z= 4x^2+4y^2;z=4

OpenStudy (anonymous):

first take 4 common frm 2nd eq then subtract both equation x n y will b canceled n z will be = to 4

OpenStudy (anonymous):

Okay.Then?

OpenStudy (anonymous):

then by comparin method find the value of x n put it in any of equ then y will automatically come

OpenStudy (anonymous):

Okay. Let's say I got the value of x, y and z. How will I calculate the volume? I am a bit rusty on triple integrals.

OpenStudy (anonymous):

hmm thts pity difficult :P

OpenStudy (anonymous):

multiply all the values didnt u get it?

OpenStudy (anonymous):

Well the process I saw online was very difficult and it required a lot more effort. Just multiplication won't help.

OpenStudy (anonymous):

ohky m sorry i dnt knw more then that m student of 1st year so0o m sorry :(

OpenStudy (anonymous):

It's alright aliza. Thanks for the insight.:)

OpenStudy (anonymous):

ur welcome nw m going just pray 4 ma tomorrows NSTC test leaving t,cr

OpenStudy (anonymous):

Sure and don't worry you will do well.

OpenStudy (anonymous):

:) inshaallah

OpenStudy (anonymous):

inshaallah.

OpenStudy (anonymous):

\[\int\limits_{0}^{2\pi}\int\limits_{1}^{2}\int\limits_{0}^{4}1 dV\]

OpenStudy (anonymous):



OpenStudy (anonymous):

that would be the projection on the x-y plane?

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

so now try elimination one of x/y and write it interms of z and the other var

OpenStudy (anonymous):

i mean y=root(z-x**2)

OpenStudy (anonymous):

 i hope u know how to find the area of the shaded part? i mean in 2-d? extend it to 3-d adding a dz

OpenStudy (anonymous):

fine let's make a guess.............state ur options i m in no mood to solve it either

OpenStudy (anonymous):

2pi,8pi,6pi and 35/4

OpenStudy (anonymous):

dose are the options... what do u think wud be a good gess

OpenStudy (anonymous):

consider two circles x**2 + y**2=4 and x**2+y**2=1 what would be the area b/n them?

OpenStudy (anonymous):

radius:2 and 1

OpenStudy (anonymous):

4pi-pi? 3pi?

OpenStudy (anonymous):

so area=3

OpenStudy (anonymous):

times 4 and we get 12pi

OpenStudy (anonymous):

nope now it's gone 3d multiply it with 4 #height ok? now it's 12 pi now approximate it to be a cone now wat's the volume of a cone?

OpenStudy (anonymous):

owwww i didnt think of that..

OpenStudy (anonymous):

4/3 * 12 * pi =8*pi ?

OpenStudy (anonymous):

1/3(pi)(r^2)(h)

OpenStudy (anonymous):

now find the closest answer

OpenStudy (anonymous):

8pi

OpenStudy (anonymous):

noo sorry it's 1/3 * 12 *pi 4pi now wat's the closest?

OpenStudy (anonymous):

idk deres 2pi and 6pi and 8pi... i gess 2pi then

OpenStudy (anonymous):

yup cone obviously has a larger volume :P guess that's it @shubhanknigam

OpenStudy (anonymous):

Thank you both of you, Tushara and Avinash. You were a big help. Guess I can sleep now.